# Accounting for motion of supernova host galaxy in statistical inference from SNIa data

### v1

@@ 5 Conclusion and Discussion
Cosmology has become a precision science in the past few decades. This has been achieved by an unprecedented increase in both the quality and quantity of data—Planck CMB temperature and polarization maps, Large galaxy surveys such as SDSS, and the detection of thousands of SNIa, etc. This also means physical effects hitherto considered beyond the ability of astronomical instruments are well within reach. However, such rapid progress also requires one to constantly improve existing statistical techniques and develop new ones. This paper is an effort in that direction.
@@
The aim of the paper is two-fold. First, we develop a general Bayesian method to analyse data with errors in both independent and dependent variables (discussed in detail in section 2.1.1). Second, we employ this statistical estimator to SNIa data to test its applicability, efficacy, and relevance to the current and future SNIa and other cosmological data sets. In particular, our aim is to discern the role of peculiar velocities of host galaxies of SNIa on the determination of cosmological parameters.
In Figures 1 and 2, we apply our proposed method on the Pantheon SNIa data for two cosmological models we employ for illustration. The figures show that the current data do not require us to account for the nonzero peculiar velocity of the SNIa host galaxies. We next apply our estimators on synthetic data, which partly confirms our results with Pantheon data, for parameters of the current data (Figures 3 and 4). However, we show that the future data (e.g. Zwicky Transient Facility (ZTF)[30] and the Large Synoptic Sky Survey Telescope (LSST) [31, 32]) will be very sensitive to peculiar velocities of SNIa host galaxies, and we would need appropriate statistical tools (section 2.1.1) to extract cosmological parameters from such a data set without bias (Figures 3 and 4).
@@
One possible improvement of our work is more detailed modelling of peculiar velocities of host galaxies. It is difficult to predict the peculiar velocities of host galaxies without N-body simulations (e.g. [19] for a recent effort on the impact of peculiar velocities on Hubble’s constant measurement). This velocity would also depend on the location of host galaxies, e.g., a host galaxy in a rich cluster of galaxies as opposed to an isolated galaxy. In addition, and importantly, velocities are correlated across 100s of Mpc (e.g. [34] for details and further references), and this correlated component needs to be accounted for in a more detailed discussion (e.g. [33]). Also, the aim of our paper is to treat peculiar velocities as nuisance parameters and to determine their impact on the interpretation of SNIa data. However, it is clear from the context that one might seek to infer cosmological peculiar velocities from the SNIa data. From cosmological data, peculiar flows are difficult to measure for individual objects and their presence is revealed through redshift-space distortion of galaxy correlation functions (e.g. [34] for details) and kSZ effect of clusters (e.g. [17, 18] for further details). Our proposed method might complement this ongoing effort to measure cosmological peculiar flows.
@@
In future work, our aim is to extend our work with a better modeling of peculiar velocities along with studying the feasibility of their detection using the SNIa data. In addition, a combined analysis using SNIa data sets Pantheon$+$, DESY5, and Union3 might improve our results, which we hope to investigate in a later work. We also expect that our work will find general applicability in the analysis of other cosmological data sets, which we might study in upcoming works.
@@ Acknowledgments
UU would like to acknowledge Yashi Tiwari for her valuable input throughout the project, particularly her insightful feedback on improving the efficiency of our Python code. UU also thanks Somnath Bharadwaj, Purba Mukherjee, and Avinash Paladi for their helpful comments.
@@ Appendix A Linear Approximation
The problem of fitting errors-in-variable models is quite straightforward for linear models. This is because the linear model allows us to marginalize the likelihood over the true regressor analytically, resulting in a very simple likelihood that depends only on the observed values of the variables [25]. Since our model is not linear, we cannot apply it exactly. We compare our method with the standard method applied to the linear approximation of the model. Here, we derive an expression for the likelihood of the linear approximation of our model.
@@
$m(z)=M+5\;\log\left(\frac{D_{\rm L}(z)}{\text{Mpc}}\right)+25\;,$ (A.1)
@@
$D_{\rm L}(z)=(1+z)\int_{0}^{z}\frac{c\;dz}{H(z)}.$ (A.2)
Considering errors in the measurement of both the apparent magnitude, $m$, and redshift, $z$, the likelihood can be expressed as
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z^{*}_{i})}{\sigma_{m_{i}}}\right)^{2}-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{z_{i}-z^{*}_{i}}{\sigma_{z_{i}}}\right)^{2}$ (A.3)
@@
$m^{\prime}(z)=\frac{5}{\ln 10}\left[\frac{1}{1+z}+\frac{1}{H(z)\int_{0}^{z}\frac{dz^{*}}{H(z^{*})}}\right]$ (A.5)
With the above approximation, eq. (A.4), the likelihood takes the form,
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z_{i})}{\sigma_{m_{i}}}\right)^{2}-\\ \frac{1}{2}\sum_{i=1}^{N}\left[\left(\frac{m^{\prime 2}(z_{i})}{\sigma^{2}_{m_{i}}}+\frac{1}{\sigma^{2}_{z_{i}}}\right)(z^{*}_{i}-z_{i})^{2}-\frac{m^{\prime}(m_{i}-m(z_{i}))}{\sigma^{2}_{m_{i}}}(z^{*}_{i}-z_{i})\right]$
@@
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\frac{\left(m_{i}-m(z_{i})\right)^{2}}{\sigma^{2}_{m_{i}}+m^{\prime 2}(z_{i})\;\sigma^{2}_{z_{i}}}$ (A.6)
Now, the likelihood depends only on the observed data. If we do not know the values of $\sigma_{z_{i}}$, as in our case, we can fix it at some value or make it a parameter. In this work, we chose to fix it at $\sigma_{z}=0.003\;$, which corresponds to the peculiar velocity of $1000$ km/s.

### v2

@@ V Conclusion and Discussion
Cosmology has become a precision science in the last few decades. This has been achieved by an unprecedented increase in the quality and quantity of data, namely, Planck CMB temperature and polarization maps, large galaxy surveys such as SDSS, the detection of thousands of SNIa, etc. This also means that physical effects hitherto considered beyond the ability of astronomical instruments are well within reach. However, such rapid progress also requires one to continually improve existing statistical techniques and develop new ones. This paper is an effort in that direction.
@@
In this paper, we have developed general and approximate Bayesian estimators (discussed in detail in Section II.1.1) to analyze data with errors in both independent and dependent variables and applied them to existing and synthetic SNIa data to test their applicability, efficacy, and relevance to current and future SNIa and other cosmological data sets. In particular, we have examined the role of peculiar velocities of host galaxies of SNIa in the determination of cosmological parameters.
We have applied our method to the Pantheon SNIa data focusing on two cosmological models. Our results show that measurement of cosmological parameters with the current data does not change significantly by including peculiar velocities of the SNIa host galaxies. However, our results become more consistent with $\Lambda$CDM when peculiar velocities are taken into account, which is significant because of implications for the nature of dark energy.
Application of our estimators to synthetic data with errors similar to the Pantheon data shows trends consistent with those observed in the results for the Pantheon data, suggesting that our assumptions about peculiar velocities are valid. We also showed that the future data with smaller magnitude errors (e.g. Zwicky Transient Facility (ZTF)Graham et al. (2019) and the Large Synoptic Sky Survey Telescope (LSST) Abell et al. (2009); Ivezić et al. (2019)) will be much more sensitive to peculiar velocities of host galaxies of SNIa; therefore, the statistical tools developed in this paper will become essential for extracting cosmological parameters from future data sets.
Our approach differs from previous studies based on N-body simulations and galaxy data (e.g., Gavas et al. (2024); Wu and Huterer (2017); Wojtak et al. (2014); Peterson et al. (2024); Davis et al. (2011); Peterson et al. (2022); Giani et al. (2024)) to assess the impact of peculiar velocities, as it is based only on the existence of the magnitude–redshift relationship, and treats the effect of peculiar velocities as noise in the measurement of supernova redshift. Our method does not need to make prior assumptions about the correlation of velocities, e.g., the correlated large-scale component and the random small-scale component as expected in the $\Lambda$CDM model. One possible extension of our work is a more detailed modeling of the peculiar velocities of host galaxies.
@@
It is difficult to predict the peculiar velocities of host galaxies without N-body simulations (e.g. Gavas et al. (2024) for a recent effort on the impact of peculiar velocities on Hubble constant measurement). This velocity would also depend on the location of host galaxies, e.g. a host galaxy in a rich cluster of galaxies as opposed to an isolated galaxy. In addition and importantly, velocities are correlated across hundreds of Mpc (see, e.g., Dodelson and Schmidt (2020) for details and further references), and this correlated component needs to be accounted for in a more detailed discussion (e.g., Garcia et al. (2020)). In addition, the aim of our paper is to treat peculiar velocities as nuisance parameters and to determine their impact on the interpretation of the SNIa data. However, it is clear from the context that one might seek to infer cosmological peculiar velocities from the SNIa data. From cosmological data, peculiar flows are difficult to measure for individual objects and their presence is revealed through redshift-space distortion of galaxy correlation functions (e.g. Dodelson and Schmidt (2020) for details) and kSZ effect of clusters (e.g. Aghanim et al. (2018); Tanimura et al. (2021) for further details). Our proposed method complements this ongoing effort to measure peculiar cosmological flows.
@@
In future work, our aim is to extend our work with a better modeling of peculiar velocities along with studying the feasibility of their detection using the SNIa data. In addition, a combined analysis using SNIa data sets Pantheon$+$, DESY5 and Union3 might improve our results, which we hope to investigate in a later work. We also expect that our work will find general applicability in the analysis of other cosmological data sets, which we might study in upcoming works.
@@ Acknowledgments
UU would like to acknowledge Yashi Tiwari for her valuable input throughout the project, particularly her insightful feedback on improving the efficiency of our Python code. UU also thanks Somnath Bharadwaj, Purba Mukherjee, and Avinash Paladi for their helpful comments. We acknowledge the use of the HPC cluster of the Raman Research Institute, Bangalore, India.
@@ Appendix A Linear Approximation
The problem of fitting errors-in-variable models is quite straightforward for linear models. This is because the linear model allows us to marginalize the likelihood over the true regressor analytically, resulting in a very simple likelihood that depends only on the observed values of the variables Press et al. (1992). Since our model is not linear, we cannot apply it exactly. We compare our method with the standard method applied to the linear approximation of the model. Here, we derive an expression for the likelihood of the linear approximation of our model.
@@
$m(z)=M+5\log\left(\frac{D_{\rm L}(z)}{\text{Mpc}}\right)+25\;,$ (19)
@@
$D_{\rm L}(z)=(1+z)\int_{0}^{z}\frac{c\;dz}{H(z)}.$ (20)
Considering errors in the measurement of both the apparent magnitude, $m$, and redshift, $z$, and assuming the errors to be Gaussian distributed, the likelihood can be expressed as
$\ln\mathcal{L}\propto-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z^{*}_{i})}{\sigma_{m_{i}}}\right)^{2}-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{z_{i}-z^{*}_{i}}{\sigma_{z_{i}}}\right)^{2}$ (21)
@@
$m^{\prime}(z)=\frac{5}{\ln 10}\left[\frac{1}{1+z}+\frac{1}{H(z)\int_{0}^{z}\frac{dz^{*}}{H(z^{*})}}\right]$ (23)
With the above approximation, Eq. (22), the likelihood takes the form,
$\ln\mathcal{L}\propto-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z_{i})}{\sigma_{m_{i}}}\right)^{2}-\\ \frac{1}{2}\sum_{i=1}^{N}\left[\left(\frac{m^{\prime 2}(z_{i})}{\sigma^{2}_{m_{i}}}+\frac{1}{\sigma^{2}_{z_{i}}}\right)(z^{*}_{i}-z_{i})^{2}-\frac{m^{\prime}(m_{i}-m(z_{i}))}{\sigma^{2}_{m_{i}}}(z^{*}_{i}-z_{i})\right]$
@@
$\ln\mathcal{L}\propto-\frac{1}{2}\sum_{i=1}^{N}\frac{\left(m_{i}-m(z_{i})\right)^{2}}{\sigma^{2}_{m_{i}}+m^{\prime 2}(z_{i})\;\sigma^{2}_{z_{i}}}$ (24)
Now, the likelihood depends only on the observed data. If we do not know the values of $\sigma_{z_{i}}$, as in our case, we can fix it at some value or make it a parameter. In this work, we chose to fix it at $\sigma_{z}=(\sigma_{v_{p}}/c)(1+z)\;$, with $\sigma_{v_{p}}=1000$ km/s.